Shift summary for eve 9/25/99

Jason Hamm (jhamm@physics.arizona.edu)
Sat, 25 Sep 1999 22:15:50 -0700 (MST)


Shift Summary for eve shift, 09/25/99

On Shift: A. Erwin, J. Hamm

Hours of beam available at all intensities: ~2.5
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 1
Hours of special runs (laser scans, muon runs...): .25

Accounting of lost time: Recovery from nm2eu power supply failure,
problems switching nm2sr polarity, rf errors, finishing up hcc gain
match

========================================================================
Runs ended this shift:

14726 - Muon
14730 - Beam_799

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Finishing HCC gain matching, muon run, waiting for
beam.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Still recovering from nm2eu supply failure.

Meanwhile, ST and JCH work on HCC gain matching
in order to counter the drift in thresholds
caused by radiation damage. We become scared when
the PMT gain drift plot looks like a random
distribution - after a few hours, we realize that
MF 9 and 11 restored incorrectly when the new CSI
HV's were installed.

After this was fixed, two outliers on the plot
remained: 1392 (+15%) and 1533 (-10%). HV's were
correct, so we decided to install new gain
constants for them and deal with the threshold
consequences later. New gain constants valid from
run 14725 onward.

19:15 Beam returns - we take muon run 14726. Afterwards
we are unable to flip polarity of nm2sr, so we
leave it as "positive" for now.

20:10 Onsplit requests UPMU01 on ktev3

20:55 After starting Beam_799 run are unable to get any
triggers through (many pipeline errors). Notice
red RF button on slowdaq window. BobT reloads RF,
and we fresh start.

CSPED updated.

21:00 Start Beam_799 run 14730.

21:35 Beam gone, linac problems - no estimate.

22:25 Ops lets us know that the reversal switch for
nm2sr is now fixed. Beam returns at 6e12. Notice
unhappy faces on l3 cpu summaries - decide to stop
run and brute force l3 processes while beam is
still poor.

23:07 Start Beam_799 run 14731. Beam spotty through
rest of shift.


========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Keep an eye on l3 cpu summaries.

Also, look out for very inefficient hcc bits, now that the gain match
is finished.